alex is 2 years old and his mother is 26 years old.In how many years'time will his mother's age be three time Alex's age?
Answer by jorel1380(3719)

Let n be the amount of years before Alex's mother is 3x his age. Then:
3(2+n)=26+n
6+3n=26+n
2n=20
n=10
☺☺☺☺